OpenStudy (anonymous):

A=pir^2 solve for r. Why isn't it A/4pi all under a radical?

OpenStudy (saifoo.khan):

from where u got 4?

OpenStudy (anonymous):

typed it up wrong its A=4pir^2

OpenStudy (saifoo.khan):

Yes, correct.

OpenStudy (anonymous):

:( got it marked wrong. Alright thanks

OpenStudy (saifoo.khan):

\[r = \sqrt{\frac{A}{4 \pi}}\]

OpenStudy (anonymous):

Yeah, I guess that shouldn't have been marked wrong

OpenStudy (saifoo.khan):

might be there is some format mistake! Mathlab suck at that. >_< grrrr.

OpenStudy (anonymous):

Looks good I think

OpenStudy (saifoo.khan):

:/